simplify each expression then evaluate the expressionfor the given value of the variable
-b^2(2b+4)+b^5,b=-1
3c^2+1+(5c)^2,c=3
20-2x^2/x,x=-2
y^-3(2y/9),y=-3
1
Expert's answer
2011-12-09T09:13:35-0500
simplify each expression then evaluate the expression for the given value of the variable
1)& -b^2(2b+4)+b^5, b=-1
-b^2(2b+4)+b^5 = -(-1)^2(2(-1)+4)+(-1)^5 = -1(-2+4)+(-1) = -1(-2+4)+(-1) = -2-1 = -3.

2)& 3c^2+1+(5c)^2,c=3
3c^2+1+(5c)^2 = 3*3^2+1+(5*3)^2 = 3^3+1+15^2 = 27+1+225 = 253.

3) 20-2x^2/x,x=-2
20-2x^2/x = 20-2(-2)^2/(-2) = 20-2*4/(-2) = 20+2*4/2 = 20+4 = 24.

4)y^(-3)(2y/9),y=-3
y^(-3)(2y/9) = (-3)^(-3)(2(-3)/9) = (-3)^(-3)(-6/9) = (-1/3^3)(-6/9) = (1/3^3)(6/9) = (1/27)(6/9) = 6/(9*27) = 2/(3*27) = 2/81.